# Inflations among quantum Grothendieck rings of type A

### v2

@@ 5.3. Categorification of quantum inflations
for any $p\in\mathbb{Z}$, where we used the notation from (3.3). Then, we have the induced homomorphism $\nu_{*}\colon P\to P$ given by $\nu_{*}(\epsilon_{a})\coloneqq\epsilon_{\nu(a)}$, which restricts to $R^{+}\to R^{+}$ satisfying $\nu_{*}(\alpha_{a,b})=\alpha_{\nu(a),\nu(b)}$ for any $a,b\in\mathbb{Z}$ with $a<b$. Let us consider the collection $\{L(\nu_{*}(\alpha_{a}))\mid a\in\mathbb{Z}\}$ of simple modules in $\mathscr{A}$, and apply the construction of Kashiwara–Park [17] to the duality datum formed by the affinizations of these simple modules. Note that the degree of the normalized $R$-matrix between $L(\nu_{*}(\alpha_{a}))$ and $L(\nu_{*}(\alpha_{b}))$ is equal to $-(\nu_{*}(\alpha_{a}),\nu_{*}(\alpha_{b}))=-(\alpha_{a},\alpha_{b})$ as computed in [13, Proposition 4.3.(vii)]. The duality datum gives rise to a $\mathbb{Z}$-graded exact monoidal functor
@@
By the universality of the localizations, it is easy to see that there exists a unique $\mathbb{Z}$-graded exact monoidal functor
@@
[…]
@@
[…]
The composition $F_{\tilde{n}}\circ F_{\nu}\colon\mathscr{T}_{n}\to\mathscr{C}_{\tilde{n}}$ can be constructed more directly following the recipe of [15, §6.1], where it is also proved that the resulting functor always respects the simple classes [15, Theorem 6.9]. It gives an alternative proof of Corollary 4.9 for $\Psi_{{\tilde{\xi}}_{0},\nu,\xi_{0}}|_{t=1}$ (and hence an alternative proof of Conjecture 2.1) without using the quantum Grothendieck rings.
@@
[…]
@@ Acknowledgments
This work was supported by JSPS KAKENHI Grant No. JP23K12955.

### v3

@@ 5.3. Categorification of quantum inflations
for any $p\in\mathbb{Z}$, where we used the notation from (3.3). Then, we have the induced homomorphism $\nu_{*}\colon P\to P$ given by $\nu_{*}(\epsilon_{a})\coloneqq\epsilon_{\nu(a)}$, which restricts to $R^{+}\to R^{+}$ satisfying $\nu_{*}(\alpha_{a,b})=\alpha_{\nu(a),\nu(b)}$ for any $a,b\in\mathbb{Z}$ with $a<b$. Let us consider the collection $\{L(\nu_{*}(\alpha_{a}))\mid a\in\mathbb{Z}\}$ of simple modules in $\mathscr{A}$, and apply the construction of Kashiwara–Park [17] to the duality datum formed by the affinizations of these simple modules. Note that the degree of the normalized $R$-matrix between $L(\nu_{*}(\alpha_{a}))$ and $L(\nu_{*}(\alpha_{b}))$ is equal to $(\nu_{*}(\alpha_{a}),\nu_{*}(\alpha_{b}))=-(\alpha_{a},\alpha_{b})$ as computed in [13, Proposition 4.3.(vii)]. The duality datum gives rise to a $\mathbb{Z}$-graded exact monoidal functor
@@
By the universality of the localizations (see the paragraph before Theorem 5.1), there exists a unique $\mathbb{Z}$-graded exact monoidal functor
@@
(5.8) ${K(\mathscr{T}_{n})\otimes_{\mathbb{Z}[t^{\pm 1}]}\mathbb{Z}[t^{\pm 1/2}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{n}}\scriptstyle{\sim}$$\scriptstyle{[F_{\nu}]\otimes 1}$${K_{t}(\mathscr{C}_{n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Psi_{{\tilde{\xi}}_{0},\nu,\xi_{0}}}$${K(\mathscr{T}_{\tilde{n}})\otimes_{\mathbb{Z}[t^{\pm 1}]}\mathbb{Z}[t^{\pm 1/2}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{\tilde{n}}}\scriptstyle{\sim}$${K_{t}(\mathscr{C}_{\tilde{n}}).}$
@@
${K(\mathscr{T}_{n})/(t-1)K(\mathscr{T}_{n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{[F_{n}]}\scriptstyle{\sim}$$\scriptstyle{[F_{\nu}]}$${K(\mathscr{C}_{n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Psi_{{\tilde{\xi}}_{0},\nu,\xi_{0}}|_{t=1}}$${K(\mathscr{T}_{\tilde{n}})/(t-1)K(\mathscr{T}_{\tilde{n}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{[F_{\tilde{n}}]}\scriptstyle{\sim}$${K(\mathscr{C}_{\tilde{n}}).}$
The composition $F_{\tilde{n}}\circ F_{\nu}\colon\mathscr{T}_{n}\to\mathscr{C}_{\tilde{n}}$ can be constructed more directly following the recipe of [15, §6.1] based on the generalized quantum affine Schur–Weyl duality in the sense of [13]. In [15, Theorem 6.9], it is also proved that the resulting functor always respects the simple classes. In this way, one may obtain an alternative proof of Conjecture 2.1 without using the quantum Grothendieck rings.
@@
${K(\mathscr{T}_{n})\otimes_{\mathbb{Z}[t^{\pm 1}]}\mathbb{Z}[t^{\pm 1/2}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{n}}\scriptstyle{\sim}$$\scriptstyle{[F_{{\tilde{\xi}},\nu,{\tilde{\xi}}}]\otimes 1}$${K_{t}(\mathscr{C}_{n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Psi_{{\tilde{\xi}},\nu,\xi}}$${K(\mathscr{T}_{\tilde{n}})\otimes_{\mathbb{Z}[t^{\pm 1}]}\mathbb{Z}[t^{\pm 1/2}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{\tilde{n}}}\scriptstyle{\sim}$${K_{t}(\mathscr{C}_{\tilde{n}}).}$
In particular, $F_{{\tilde{\xi}},\nu,\xi}$ sends simple objects of $\mathscr{T}_{n}$ to simple objects of $\mathscr{T}_{\tilde{n}}$.
@@ Acknowledgments
The author would like to thank the anonymous referees for their useful comments and suggestions. This work was supported by JSPS KAKENHI Grant No. JP23K12955.
Declarations
The author has no relevant financial or non-financial interests to disclose. We do not analyze or generate any datasets, because our work proceeds within a mathematical approach.